Personal tools
  • We're Hiring!

You are here: Home Community Minutes Conference Calls 2010 2010-04-08 Thursday Meeting

2010-04-08 Thursday Meeting

Attending: Brian, Chris, Andrew, Will, Jean-Marie, Ola, Josh, Melissa, Carlos, Jason


Remember: Agenda must be complete (with estimated times) on the day before the meeting. Any additions after that must go at the bottom (AOCB)

  1. Accepting minutes from last meeting

  2. Matters Arising (<10 mins)

    • Everyone: please finish updating your shoola trac email information!!

    • Will, J-M, Donald: meet tomorrow about RIO changes needed

    • Donald: Email David to find out his requirements and dead line for FLIM

    • Donald: split #1985 into refined tickets

    • Andrew: Need to add tickets for changes from last meeting

    • Andrew: Add Brian Jones to the Dev Contact List

    • Colin: Add tickets for unit testing

    • Chris: Add your additional tickets for bio-formats to sprint 6

    • Will: Need to have another discussion with Christof & Ingvar

    • Will: Find out if Christof is coming mid-April

    • Will: Send update/screenshots/show where we're at to Paul, Steve L., and Christof

    • Everyone: Next Tuesday organise going through all tickets and assign to sprints

    • Chris: Answer the question on the forums about atomic SPW imports

    • Chris: Answer the guy about var log question

    • Josh/Chris: Make the settings changes for permissions ASAP

    • Melissa/Will: Discuss EM/SPM/SEM metadata

    • Melissa: Estimate work effort for 4.2 next week

    • Chris/Melissa/Curtis: discuss the use of OME Model objects

    • Chris/Melissa: add bioformats version to imported properties

    • Andrew, J-M, Ola: metadata diagram making public

  3. Sprint - Status Update (30 mins)

  4. Hudson builds for windows (2 mins)

  5. Paris Meeting (5 mins)

  6. Email and Forum Review - checked at 13:10 (5 mins)

    • Email - Arcieri, Kenneth - Connection lost exception
    • Forums - Can't create movie link, Apache mod_python failure after reboot link, Convert to OME-TIFF link,
    • QA - Importer bug (Ice.UnknownLocalException), Importer bug (aperio file version), Importer Bug (first upload), Importer Bug (Unable to locate Pixels), Importer Bug (Ice.SocketException), Web (Security Violation), Web (unable to open database)
  7. AOCB (<5 mins)


  • Last Meetings Minutes: Approved

  • Action Items from Last Meeting

    • Did everyone review their tickets in preparation for next week? No.
  • Sprint status

    • Would be good to
    • What does the graph say?
    • Chris: Melissa is ahead :) Largely on track
      • testing of metadata levels is in (with more logging)
    • Jean-Marie: need to close the permission chapter. Important!
    • Need to make sure that our clients are properly working.
    • FS will be a proof-of-concept (re-import blocker)
    • Jean-Marie: realistically, large number of things on roadmap
      • Need something solid for permissions, bioformats, scripting, web
      • behind on the scripting front. need to push on FLIM
      • Need to deliver something for FLIM this iteration
    • For next week
      • Full-court press permissions testing (Monday) Discuss Tuesday.
        • Jason: drawing a line under it on Tuesday?
        • Chris: if changes are sweeping may need more time
      • FS w/ Colin: reimport of metadata, ...
      • Model changes: SPW, new filter, ...
      • Smaller scripting discussions
        • Will upgrades his scripts with new features
        • Testing the full stack
        • Scripts running on mage so everyone can try them * Blocked right now by permissions (#2073) * Need to discuss early in the week
        • Multiple people running them at the same time
      • Donald on FLIM
        • Database change from Donald needed for this
      • @Andrew give us a visual summery of 'image down' model after changes (filter, etc)
    • Will: working on scripts. Needs enumerations, min/max, etc.
    • Testing: is everyone aware of where/how? (See #1736 HUDSON GREEN ticket)
    • Holding off Ice changes for now in favor of db changes?
      • Don't want to delay this too long
      • Jean-Marie: upgrade of Ice for OMERO 4.2.1. Lets see nearer to release
  • Builds

    • gre-mporter;
    • 4.1 we were IO bound
    • Chris: has put a request in for fiber channel disk for hudson. When? Unknown.
    • May need dedicated disks for Windows v. Virtual images?
  • Paris

    • Poster session
    • How many people? Trickling & the unregistered folk make it difficult
    • Will: workshops are consecutive
    • Should open up the walls between rooms? (Was cramped last year)
  • Forums/emails/QAs

    • Chris: jabber notifications don't work
    • Melissa will reply on OME-TIFF after fixing the bug
    • Web: Ola needs more info.
    • During testing, easier to create a ticket rather than use QA.
    • @Best practice: everyone on the Dev Team use tickets
    • @Best practice: put stack traces in comments not description
  • AOCB

    • Donald. 21st April. Presentation. @Andrew will provide poster to Donald
      • Jason: talks are better
      • refactor poster to better user space

Action Items

  • Jean-Marie: Please post all presentations to the plone
  • Ola/Carlos/Curtis: Add email details to Shoola Trac
  • Everyone: move tickets that can't be done by 4.2 to 'future' (MUST DO BY NEXT THURSDAY)
  • Andrew: give us a visual summery of 'image down' model after changes (filter, etc)
  • Josh/Chris: Do a diff of the DSL / compile and see what classes break (Do Friday)
  • Everyone: Review Jason's email with preliminary Paris meeting agenda
  • Everyone: Re-subscribe to the forums to receive email notifications
  • Andrew: Get a poster together for Donald's talk on the 21st

Create Tickets

  • Chris: Add bioformats version to imported properties

Best Practices

  • Everyone on the Dev Team use tickets
  • Put stack traces in comments not description
Document Actions